OpenStudy (anonymous):

Convergent or divergent 16-8+4-...

OpenStudy (freckles):

do you know what kinda series that is?

OpenStudy (freckles):

hint: is there a common ratio?

OpenStudy (freckles):

if so what is it?

OpenStudy (anonymous):

No, I have to figure out the infinite geometric series

OpenStudy (freckles):

there is a common ration since it is a geometric series

OpenStudy (freckles):

the geometric series here has a ratio of (-1/2)

OpenStudy (freckles):

do you see that now?

OpenStudy (freckles):

has a common ratio of (-1/2)*

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So is it divergent?

OpenStudy (freckles):

now since you just need to determine if it is divergent or convergent you just need to tell me if you have |r|<1 --- if |r|<1 then it is convergent if |r|>=1 then it is divergent

OpenStudy (freckles):

you have r=-1/2 so which inequality is true ?

OpenStudy (anonymous):

oh, its convergent. <

OpenStudy (freckles):

yep yep

OpenStudy (anonymous):

thank you! can you answer another?

OpenStudy (freckles):

okay

OpenStudy (anonymous):

Find the sum of each infinite series, if it exists 440+220+110+...

OpenStudy (freckles):

let's see what kinda series is this?

OpenStudy (freckles):

can you tell?

OpenStudy (freckles):

have you looked for a common ratio?

OpenStudy (anonymous):

i think its an infinite one

OpenStudy (freckles):

if there is a common ratio, it is a geometric series

OpenStudy (anonymous):

i looked for one but really couldnt figure it out

OpenStudy (freckles):

220/440=? 110/220=?

OpenStudy (anonymous):

okay so 1/2

OpenStudy (freckles):

yes both of them are equal to 1/2 therefore there is a common ratio of 1/2 therefore since it has a common ratio it is a geometric series

OpenStudy (freckles):

use the same test as we did before

OpenStudy (anonymous):

you mean the convergent and divergent thing?

OpenStudy (anonymous):

if so its convergent since its less than 1

OpenStudy (freckles):

the geometric convergeny test: if |r|<1 then it is convergent if |r|>=1 then it is divergent

OpenStudy (freckles):

and yes

OpenStudy (freckles):

if we had 2,4,8,... this would be a different story

OpenStudy (freckles):

the common ratio here would be 2

OpenStudy (freckles):

so 2+4+8+... would diverge

OpenStudy (anonymous):

and divergent since its over 1?

OpenStudy (freckles):

right

OpenStudy (anonymous):

okay yeah i get it.! so if i were to find the sum of the last series? how would i do that?

OpenStudy (freckles):

if we have -2,4,-8,... then the common ratio is -2 and so -2+4+(-8)+... is divergent since |-2|=2>1

OpenStudy (freckles):

the formula is this: \[\sum_{n=0}^{\infty}a r^n=\frac{a}{1-r}\]

OpenStudy (freckles):

so you have a=440 and r=1/2

OpenStudy (freckles):

enter those in

OpenStudy (anonymous):

hmmm cant seem to find it on my calculator. D:

OpenStudy (freckles):

find what/

OpenStudy (anonymous):

the formula to plug the numbers into

OpenStudy (freckles):

\[\frac{440}{1-\frac{1}{2}} =?\] a division sign and subtraction sign?

OpenStudy (freckles):

(440) div (1-1 div 2)

OpenStudy (freckles):

so you think your calculator has the formula in it?

OpenStudy (freckles):

what kinda calculator do you have?

OpenStudy (anonymous):

TI 84 Plus by Texas Instruments

OpenStudy (freckles):

it looks like you can use the calculator to by pass using the formula

OpenStudy (freckles):

but you have to know the put it in sigma notation

OpenStudy (freckles):

440,220,110,... first number is 440 440,440(1/2),440(1/2)^2,... this can be written as \[\sum_{k=0}^{\infty}440 \cdot (\frac{1}{2})^k\] so it says you can enter seq(440(1/2)^k, 0,inf) it looks like

OpenStudy (freckles):

then you can use the sum command

OpenStudy (freckles):

sum(Ans)

OpenStudy (freckles):

that seems like a lot of work

OpenStudy (freckles):

I think using the a/(1-r) is pretty easy

OpenStudy (anonymous):

okay yeah that way seemed a bit easier

OpenStudy (anonymous):

so does that big greek looking E have any significance?

OpenStudy (freckles):

is that what you are confused about?

OpenStudy (freckles):

that big sigma thing means we are adding a whole bunch of terms

OpenStudy (freckles):

example: \[\sum_{k=0}^{5}k^2 \text{ means } 0^2+1^2+2^2+3^2+4^2+5^2\]

OpenStudy (freckles):

the sigma sign tells us we are adding the k=0 underneath means we going to starting adding terms starting with k=0 the 5 on top means that is where we will stop

OpenStudy (anonymous):

well i missed class today. Then learned a new section and when i got my work after school she gave me a 40 problem packet and told me it was due tomorrow. so im just now trying to learn what all these words and symbols mean.

OpenStudy (freckles):

so instead of writting the series in sigma notation I could have wrote: \[a+ar+ar^2+ar^3+ar^4 + \cdots =\frac{a}{1-r}\]

OpenStudy (freckles):

where |r|<1

OpenStudy (anonymous):

so to do these it has to be convergent.?

OpenStudy (freckles):

yes you can only use that formula if you have |r|<1

OpenStudy (anonymous):

ok sounds good

OpenStudy (anonymous):

ok so when i plug in the numbers i got 880. is that the final answer?

OpenStudy (anonymous):

well if youre still there. you were a big help thank you and sorry im a dumbass

OpenStudy (freckles):

880 is correct

OpenStudy (freckles):

my notifications stopped working